OpenStudy (anonymous):

Tom’s brother saw this advertisement for a special at Muscle Fitness. He tells Tom that he feels it is a better deal than Platinum Gym. Muscle Fitness join this month and save $20 off your registration feel! membership is only $15 per month! Regular Registration Fee: $80 (one time fee) Additional Fee: $5 per month equipment maintenance (mandatory) (a) Write an expression to represent the membership cost of Muscle Fitness for x months. (b) Is Tom’s brother right? Explain why or why not.

OpenStudy (anonymous):

a = 80-20+15*x=? x = months

OpenStudy (anonymous):

ok here a = 80-20+15*x+5*x x= number of months b i dont know because you only show for the one gym